OpenStudy (anonymous):

any one know what are the observable values of momentum and how are they calculated

OpenStudy (anonymous):

@blacklotus ?

OpenStudy (anonymous):

it is the EXPECTATION VALUE in a state\[\Psi\] , one means OBSERVABLE VALUE of momentum operator. This means the following: \[\Large\langle \hat{p} \rangle = \int \int \int \Psi^*\left( -i\hbar\overrightarrow{{\nabla\Psi}}( r)\right)\; dxdydz \]

OpenStudy (anonymous):

\[\Large \overrightarrow{\nabla\Psi} =\bf (\partial_x\Psi,\, \partial_y\Psi,\, \partial_z\Psi) \]